OpenStudy (anonymous):

Is there a formula for 3 (or more)-digit numbers similar to this? ab*cd=xyz z=b*d y=a*d+b*c x=a*c

OpenStudy (anonymous):

@Michalas @amistre64

OpenStudy (amistre64):

not sure what the question is asking

ganeshie8 (ganeshie8):

what would you do if b*d results in two digits ?

ganeshie8 (ganeshie8):

i think the OP is looking for a shortcut for multiplying two 3 digit numbers...

ganeshie8 (ganeshie8):

for example : 28 * 37 b*d = 8*7 = 56

OpenStudy (amistre64):

ab cd --- ac + bc ad + bd ----------- x y z so its just multiplication ....

OpenStudy (amistre64):

you could develop one ... just multiply 3 digit numbers

OpenStudy (amistre64):

abc xyz ----- ax + bx + cx ay + by + cy az + bz + cz --------------------- m n o p q

ganeshie8 (ganeshie8):

looks like we get a formula similar to determinant of 3x3 matrix to me :)

OpenStudy (amistre64):

what? thats just crazy talk .... absolutely mad i tell ya

ganeshie8 (ganeshie8):